what is 3 radical 5 multiplied by 4 radical 5
Answer by Boreal(15235)

3 sqrt (5)*4 sqrt (5)
you can multiply in any order you wish.
3*4*sqrt(5)*sqrt (5)=
3*4*5, since sqrt (5)*sqrt (5)=5
3*4*5=60